OpenStudy (anonymous):

Find all the zeroes for the function: f(x) = 3x^4 - 6x^2

OpenStudy (anonymous):

take out 3x^2 so: \[3^{2}(x ^{2}-2) = 0\] \[3x ^{2}(x-\sqrt{2})(x+\sqrt{2}) = 0\] \[x = 0,\sqrt{2},-\sqrt{2}\]

OpenStudy (anonymous):

3^2 in top line is supposed to be 3x^2, however it does not change the answer.